Nathan has two brothers, Conner and Josh. If we add 2 to Conner's age, it is exactly 1/2 the age of their older brother Josh, who is 12. How old is Conner?
I got the answer of 4 years old.
Now here is the problem. How many years will it be until Conner's age is 1/2 of Josh's age?
I know it is 4 years. But I figured it out through trial and error and need to know the equation and how to set it up. Thanks.

Answer by stanbon(75887)

Nathan has two brothers, Conner and Josh.
If we add 2 to Conner's age, it is exactly 1/2 the age of their older brother Josh, who is 12. How old is Conner?
---
Equation:
(c+2) = (1/2)(12)
c+2 = 6
c = 4
========
I got the answer of 4 years old.
OK
================================================
Now here is the problem. How many years will it be until Conner's age is 1/2 of Josh's age?
Equation:
(c+x) = (1/2)(12+x)
4+x = 6 + x/2
(1/2)x = 2
x = 4 yrs
--------
Cheers,
Stan H.
============

I know it is 4 years. But I figured it out through trial and error and need to know the equation and how to set it up.